OpenStudy (anonymous):

32.5% of what is 12.48?

OpenStudy (anonymous):

Medal plz What is 32.5 percent of 12.48 - step by step solution Simple and best practice solution for 32.5% of 12.48. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the calculator fields your own values, and You will get the solution. What is % of ? is what percent of ? To get the solution, we are looking for, we need to point out what we know. 1. We assume, that the number 12.48 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 12.48 is 100%, so we can write it down as 12.48=100%. 4. We know, that x is 32.5% of the output value, so we can write it down as x=32.5%. 5. Now we have two simple equations: 1) 12.48=100% 2) x=32.5% where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 12.48/x=100%/32.5% 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for what is 32.5% of 12.48 12.48/x=100/32.5 (12.48/x)*x=(100/32.5)*x - we multiply both sides of the equation by x 12.48=3.07692307692*x - we divide both sides of the equation by (3.07692307692) to get x 12.48/3.07692307692=x 4.056=x x=4.056

OpenStudy (abb0t):

Nice copy paste.

OpenStudy (texaschic101):

you need to learn how to do this fireball, instead of just getting answers. You have asked several questions that have to deal with percents....try to learn how to do this so you won't be asking the same sort of questions

OpenStudy (abb0t):

You tell em, texas!!! WOOHOO!!

OpenStudy (anonymous):

Suck my retriceabbot you fluttering owletgot

OpenStudy (anonymous):

You stupid f*uck

OpenStudy (abb0t):

YEAH, KEEP DREAMING!! GO DRINK UR JUICY-JUICE AND GO NAP-NAP, KID!! And LEARN UR BASIC MATH!